OpenStudy (anonymous):

what part of a yard is one-half of a foot? A. 1/12 B. 1/4 C. 1/3 D. 1/6 E. 1/2

OpenStudy (anonymous):

A foot is 1/3 of a yard... so half a foot is half of 1/3, which is 1/6. You have to multiply the two fractions.

OpenStudy (debbieg):

A foot = 1/3 of a yard. so 1/2 of a foot = (1/2)(1/3) of a yard, right? So just take the product.....

OpenStudy (anonymous):

D